G. Kutateladze

N. Muskhelishvili Institute of Computational Mathematics, Georgian Academy of Sciences Tbilisi, Georgia

ON RIEMANN–HILBERT PROBLEM FOR PLANE WITH CUTS

An explicit solution to Riemann-Hilbert problem for plane with two cuts is given.

Two cases are considered: the case of two straight cuts and the case with an arc-wise and straight cuts with a certain symmetry with respect to each other.

The solutions are constructed in Cauchy type integrals.